OpenStudy (anonymous):

A non scale model of the earth has a diameter of 12 in. Within an outer clay layer, there is a core made of a golf ball diameter of 2 in. What is the volume of the clay? Use 3.14 to approximate pi and round your answer to one decimal place.

OpenStudy (anonymous):

The area of a cylinder is A=2πrh+2πr2 Basically, you want to find the Area of the model Earth minus the area of the gold ball. Remember, in the equation of above, r mean radius, which is half the diameter.

OpenStudy (anonymous):

Equation messed up, \[A = 2\pi rh+2\pi r ^{2}\]

OpenStudy (anonymous):

Did you figure it out? Or do you need further explanation?

OpenStudy (anonymous):

no I didn't

OpenStudy (anonymous):

The shape of the earth is a sphere. Since the model is composed of two layers, it means that the golf ball is the core and the clay forms a spherical shell around the golf ball. For a sphere of radius r \[Volume = \frac{4}{3}\pi r^3\] since we have a sheel For the inner sphere we use the formula above and then we consider the outer sphere R = complete radius of the model = 6 inches (12 inches is the diameter so half of it gives R)\[V_2 = \frac{4}{3}\pi R^3\] for the volume of the clay used subtract the volume of the golf ball from the total volume of the model. \[Volume - of -clay-used = \frac{4}{3}\pi 6^3- \frac{4}{3}\pi 2^3\]

OpenStudy (anonymous):

correction # "shell"- above the figure

OpenStudy (anonymous):

so the answer is what

OpenStudy (anonymous):

Volume of the clay used = 870.826666 cubic inches ~ 871 cubic inches

OpenStudy (anonymous):

the answer is 871